OpenStudy (anonymous):

Two sides of a right triangle are -2 square root of 5 and negative square root of 5 what's the third side

OpenStudy (dan815):

pythagorean theorem

OpenStudy (anonymous):

I've tried it but don't get the square roots

OpenStudy (anonymous):

A triangle side cannot be negative. Either you copied the question incorrectly or there is an error on whatever you copied this from.

OpenStudy (anonymous):

That's how it's written in my calculus book

OpenStudy (wolf1728):

Despite the fact that you are dealing with a triangle with negative sides, you can get an answer: Sides are -2 * sqrt(5) -sqrt(5) So (hypotenuse)² = (-2 * sqrt(5))² + (-sqrt(5(hypotenuse)² = 20 + 5 (hypotenuse)² = 25 hypotenuse = 5

OpenStudy (anonymous):

How do you know the hypotenuse isn't -5? -5 makes the pythagorean theorem true in this case also. I'm sorry but this question is just ridiculous.

OpenStudy (wolf1728):

Yes, the question does seem ridiculous especially the negative lengths. I suppose that elusive side must be -5 . Didn't Euclid say all lengths of a triangle must be all negative or all positive. (No, I din't think Euclid said that)

OpenStudy (dan815):

maybe they mean something liek this

OpenStudy (dan815):



OpenStudy (dan815):

having negative x and y compnents

OpenStudy (anonymous):

No caluculs textbook would state a problem exactly like the original poster her presented it, that's for sure.

OpenStudy (anonymous):

@dan815 it's like how you drew it

OpenStudy (anonymous):

Why didn't you say so at the beginning? That triangle does not have negative side lengths. Just because the edges go into the negative parts of the XY axis doesn't mean the lengths are negative.

OpenStudy (anonymous):

That's how it is written in my paper

OpenStudy (wolf1728):

Congratulations to dan815 for figuring out what the question should be.

OpenStudy (anonymous):

So how do I get the hypotenuse